OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

OpenStudy (anonymous):

-2(x + 3) = -2(x + 1) - 4 -2x - 6 = -2x - 1 - 4 -2x + 2x = -1 - 4 + 6 0 = -5 + 6 0 = 1 (no solution)

OpenStudy (espex):

There was a math error, that is not the solution to this problem.

OpenStudy (anonymous):

oops....let me do it again... -2(x + 3) = -2(x + 1) - 4 -2x - 6 = -2x - 2 - 4 -2x + 2x = - 2 - 4 + 6 0 = 0 ANSWER: all real numbers I am sorry about that....thanks eSpex for catching my error